H3PO4 is a tribasic acid and one of it`s salt is NaH2PO4. What volume of 1M NaOH solution should be added to 12g of NaH2PO4 to convert it into Na3PO4?

Grade:11

1 Answers

Vikas TU
7 years ago
Write first the chemicl eqn. ad then use mole – concept – 1
NaH2PO4 + 2NaOH ------------> Na3PO4  +  H2O
12/120  = 0.1 mol
NaOH moles = 0.1*2 = 0.2 mol
0.2 mol =1 * V
v = 0.2 l
or 200 ml.